What Is The Molar Solubility Of Nickel(II) Sulfide In 0.091 M CN-? For Nis, Ksp = 3.0 10-19; For Ni(CN)4²-, Kf,= 1.0x 10³¹. Show All Equilibrium Equations Involved Here.




See the answer
Answer #1

Given :

For NiS, Ksp = 3.0*10-19

For [Ni(CN)4]2-, Kf = 1.0*1031

............... Ksp = 3.0*10-19

..... Kf = 1.0*1031

____________________________________________________

........Koverall = Ksp*Kf = (3.0*10-19)*(1.0*1031) = 3.0*1012

Now, Koverall = 3.0*1012 = [Ni(CN)42-] [S2-] / [CN-]4

Set up ICE table for the overall reaction to calculate the equilibrium concentrations:

4CN- NiS(s) Ni(CN)42- S2-
Initial(M) 0.091 -- 0 0
Change(M) -4s -- +s +s
Equilibrium(M) 0.091-4s -- s s

Substitute the equilibrium values in the Koverall expression.

Koverall = 3.0*1012 = s*s / (0.091-4s)4 s2 / (0.091)4 = s2 / 6.86*10-5

s2 = 6.86*10-5 * 3.0*1012 = 20.58*107 = 2.058*108

Take square root of both sides.

s = 1.4*104 M

So, the molar solubility of NiS is 1.4*104 M.